# IEEEexample.bib

### v2

@@ II-E Multi-task Learning I (MTL-01)
Considering that the loss function of Mask R-CNN measures instance segmentation and does not consider error propagation, we design a serial network (MTL-01), which takes the segmentation results of Mask R-CNN as input to ResNet-50 and calculates additional loss terms for back-propagation to assist in optimizing Mask R-CNN (Fig. 1). The ResNet-50 in the serial network is a pre-trained model, whose parameters are not updated. During the training process, the erroneous (over-)segmentation masks from Mask R-CNN are mainly categorized into two types. The first type is small false positive patches escaping from NMS(Non-Maximum Suppression) filtering[he2017mask], which were only found to slightly contribute to errors in subsequent emission rate estimation rather than main error source of the serial network. Simulation experiments indicate that the IME value of plumes is almost never below 300. Hence, we disregard false positive patches with IME values under 300 pixels, addressing this type of error using an object detection loss term. The second type of error consists of large false positive patches, often wrongly separated from overlapping plumes, causing substantial inaccuracies in emission rate estimation. To mitigate this, we introduce an additional loss term, $loss_{ER}$, for MTL-01. While $loss_{ER}$ is still based on smoothL1, segmentation results with IME values below 300 pixels are excluded from its calculation. Missed masks are treated as having an emission rate estimate of 0 kg/h, and the loss is computed accordingly:
$loss_{ER}=\begin{cases}loss_{SmoothL1}(y_{i},\hat{y_{i}})&\text{if }y_{i}\in TP\\ loss_{SmoothL1}(0,\hat{y_{i}})&\text{if }y_{i}\in FP\end{cases}$ (15)
@@
Multi-task learning I (MTL-01) is a sequential network that includes a Mask R-CNN (for plume segmentation) and a ResNet-50 (for plume emission rate estimation). The Mask R-CNN outputs segmented plumes, which are then processed by ResNet-50 in real time to compute the emission rate estimation loss. The overall loss for MTL-01 is the weighted sum of the individual losses from the two networks. ”Multiplume” refers to a methane concentration distribution map with multiple plumes, including noise. The ResNet-50 is a 50-layer residual network. FCN (Fully Convolutional Network) is a type of convolutional neural network that integrates low-level and high-level image semantic information through skip connections. ”Heads” in Mask R-CNN are responsible for object detection and mask extraction. MLP stands for multi-layer perceptron. Multi-task learning II (MTL-02) is a network designed to detect methane concentration and methane plume boundaries concurrently. It is composed of a Mask R-CNN (for plume segmentation) and a U-Net (for methane concentration inversion) sharing a ResNet-50 feature extractor. While the decoding structure of the Mask R-CNN remains the same, the U-Net’s decoding structure is adapted to include four skip connections, followed by an ASPP structure and two 1x1 convolutions. ”Simulated EnMAP” refers to EnMAP data that include added methane plume signals. ASPP consists of a 1x1 convolution layer (used for channel reduction) and several parallel atrous convolution layers. The dilation rates of these atrous convolution layers are adjustable for extracting features at various scales. ”Conv 1x1” denotes a 1x1 convolution layer, ”Dilated Conv” indicates an atrous convolution layer, and ”Pool” represents a pooling layer.
@@ II-F Multi-task Learning II (MTL-02)
The task of methane concentration inversion involves predicting pixel-level values from hyperspectral images, whereas methane plume segmentation focuses on extracting plume masks from the same type of images. Due to their similar structure and logical connection, employing multi-task learning models could potentially enhance accuracy for both methane concentration inversion and plume segmentation. Therefore, we developed a multi-task learning framework that combines U-net and Mask R-CNN architectures.
In MTL-02 (Fig. 1), the network architecture of Mask R-CNN remains unchanged, and the U-net decoding structure is connected to the mask R-CNN backbone network. Considering the output sizes of the stages in the ResNet-50 backbone network, we added an additional output to stage 0 of ResNet-50 to ensure that the size of the U-net decoding the result matches the input size, similar to LinkNet[zhou2018d]. It should be noted that the deepest output of ResNet-50 is not involved in the construction of the U-net decoding structure but only participates in the FCN[long2015fully] structure of Mask R-CNN. Furthermore, considering that there is a max-pooling layer between the two outputs of ResNet-50, which may lead to underfitting of the U-net decoding structure, we added an Atrous Spatial Pyramid Pooling (ASPP) layer composed of two dilated convolutions immediately after the U-net decoder, with dilation rates of 2 and 4.
ASPP (Atrous Spatial Pyramid Pooling)[chen2017deeplab] is based on dilated convolutions and SPP (Spatial Pyramid Pooling)[he2015spatial]. It is used to address the contradiction between reducing the feature map resolution as much as possible and capturing a large receptive field when extracting image features. ASPP uses dilated convolutions with different sampling rates to capture the context of an image at multiple scales. In ASPP, different dilated convolutions with different rates are applied to the input feature map and then merged. To enhance the model’s receptive field, the output feature maps of the dilated convolution with different sampling rates are average-pooled.
For the loss function of $MTL-02$, we adopted the classical weighted average of individual loss functions for multiple tasks[bruggemann2021exploring]:
@@
where $loss_{U\--{net}}$ is the loss function of U-net, $loss_{MaskR\,\--{CNN}}$ is the loss function of Mask R-CNN, and $w_{1}$ and $w_{2}$ are the weights for respective loss components.
To determine the weights of the MTL-02 loss function, we used the Dynamic Weighted Average (DWA) algorithm[liu2019end], the idea of which is that different tasks have different levels of difficulty and learning speed. Instead of setting different learning rates for each task, it is better to dynamically adjust the weights so that each task learns at a similar pace. In simple terms, tasks with faster loss reduction are assigned smaller weights, while tasks with slower loss reduction are assigned larger weights.
$w_{k}(t)=\frac{K\cdot e^{\frac{\lambda_{k}(t-1)}{T}}}{\sum_{i=1}^{K}e^{\frac{\lambda_{i}(t-1)}{T}}}$ (18)
$\lambda_{k}(t-1)=\frac{L_{k}(t-1)}{L_{k}(t-2)}$ (19)
where $K$ is the number of tasks, $\lambda_{k}(t-1)$ denotes the rate of loss reduction for the $k-th$ task during the $(t-1)-th$ training iteration, and $L_{k}(t-2)$ represents the loss for the $k-th$ task during the $(t-2)-th$ training iteration. When $T$ is a constant, $T=1$ indicates that $w$ is equal to the result of softmax; However, as $T$ increases significantly, $w$ approaches 1, leading to equal loss weights for all tasks.
@@ III Experiments
III-A Research areas and data
III-A1 Research areas
Situated in the southern part of China, Hong Kong is predominantly marked by hilly terrain. The region, which is highly urbanized, experiences a humid and rainy climate, with an average annual precipitation exceeding 2000 millimeters. Within Hong Kong, there are two key landfills: the Western New Territories (WENT) Landfill and the North East New Territories (NENT) Landfill. Since most of the PRISMA images were obtained from the WENT region, our generalization verification is mainly concentrated in this area. The Tuen Mun landfill (Fig. 2), which spans approximately 27 hectares, is located in the southwest of the Tuen Mun district near Shenzhen Bay. It is one of the limited waste disposal sites sanctioned by the government in the Hong Kong Special Administrative Region and primarily receives municipal solid waste from various areas, including Hong Kong Island, Kowloon, and New Territories.
@@ III-A2 Hyperspectral Data
The PRISMA satellite was launched on March 22nd, 2019, and offers hyperspectral images of global coverage with a spatial resolution of 30 m. The spectral smile is less than 5 $nm$, the spectral resolution is better than 12 $nm$ in a spectral range of 400-2500 nm (VNIR and SWIR regions) [cogliati2021prisma], and datasets are open-access. We have collected several PRISMA images of WENT, which are: $20220102$, $20220119$, $20220914$, $20230217$, $20230301$, $20230307$, $20230416$. These raw images will be used to verify the generalizability of deep learning models[thorpe2013high][foote2020fast].
@@
The study site is the WENT landfill found in the Tuen Mun District of the Hong Kong Special Administrative Region. WENT is located on a slope near Shenzhen Bay, as shown by the blue box on the map.
@@
III-B Generation of snapshots for plumes of different emission rates and wind speeds
Our simulated plumes were generated using palm software[2001PALM][maronga2015parallelized], using the average atmospheric pressure at sea level in Hong Kong (0.96 hPa) and a temperature of 296K. A specialized smog mode was implemented, including a methane dispersion feature, to simulate the diffusion of methane point sources within the troposphere from 0 to 10,000 meters. We defined four gradients for the methane point source emission rate: 500kg/h, 1000kg/h, 1500kg/h, and 2000kg/h. The wind direction was set from west to east, with wind speed gradients of 1m/s, 2m/s, 3m/s, 4m/s, 5m/s, 6m/s, 7m/s, 8m/s, 9m/s, and 10m/s (Fig. 3). For each wind speed gradient, the methane point source emission rate gradients were simulated in sequence. The simulation period covered 0 to 2.5 hours, with snapshots taken at 30-second intervals from 1 to 2.5 hours, resulting in a total of 7200 plume snapshots. Among them, the simulated plume emission data with wind speeds of $2m/s$, and $10m/s$ were used to create the validation dataset, the simulated plume emission data with wind speeds of $1m/s$, and $9m/s$ were used to create the test dataset, while the remaining plume emission data were used to create the training dataset.
@@ III-B Generation of snapshots for plumes of different emission rates and wind speeds
This figure illustrates the methane plume at various emission rates and wind speeds, modeled with the Palm software. Each image was taken 180 minutes post-emission initiation, with simulation parameters specified in Section II.B.
@@ III-C Generation of multi-task learning datasets
By taking a snapshot of a plume and averaging the methane concentrations at different heights, a 2D single plume with an emission rate of m kg/h is obtained. Multiplying all pixel values of the 2D plume by a factor of a results in a modified emission rate of m*a kg/h, which serves as the label for the sample in the emission rate dataset. To ensure the diversity of plumes, we randomly generate a threshold value between 0.05 and 0.10, and pixels in the 2D plume that are below this threshold are set to zero. Finally, a rotation enhancement of -170° to 170° is applied to the 2D plume, serving as the image counterpart to the label in the emission rate dataset sample. Plume segmentation is an instance segmentation task. To begin with, we create an empty $256\times 256$ image filled with zeros. Then, we select $N$ images from the emission rate dataset (where $N\leq 3$ and $N$ can be 0), and superimpose them randomly onto the all-zero image to obtain a multi-plume image. Simultaneously, we identify the boundaries of each plume and assign corresponding labels to the segmentation dataset. If $N=0$, the dataset will have no labels. Finally, we randomly generate $256\times 256$ Gaussian white noise and add it to the multi-plume image[jongaramrungruang2022methanet][gil2002efficient], which serves as the image for the segmentation dataset. Furthermore, the overlap ratio of the two plumes’ could not exceed $15\%$:
@@
The inversion data are derived from the multi-plume images generated during the creation of the segmentation dataset, where these multi-plume images are used as labels for the inversion dataset samples. Initially, a multi-plume image is obtained and checked for emptiness. If found empty, a random EnMAP image is selected as the sample image. Conversely, if the multi-plume image is not empty, methane transmission spectra corresponding to each pixel value are constructed column by column. For columns entirely composed of zeros, the transmission spectrum is set to a $256\times 1\times 41$ matrix with all elements equal to 1. For columns with non-zero elements, an optical depth matrix of size $256\times 1\times K$ (where $K$ is the number of elements in the absorption cross-section data) is generated using the method described in Section II.A. The negative logarithm of the pixel values in the optical depth matrix is then taken to derive the original transmittance matrix.
Initially, the central wavelengths and half-bandwidths of the 41 EnMAP bands are obtained to build the spectral response function (SRF)[green1998spectral][2008Toward]. Subsequently, the SRF is applied to perform normalized convolution along the $K$ dimension of the original transmittance matrix, producing a $256\times 1\times 41$ ’column’ transmission spectrum matrix. After each column of the non-empty multi-plume image is processed, all the transmission spectrum matrices are concatenated in their original column sequence to create a $256\times 256\times 41$ transmission spectrum matrix.
Finally, the transmission spectrum matrix is combined with random EnMAP data via the Hadamard product to generate the image for the non-empty multi-plume. It is crucial to highlight that, because the simulated methane plumes are concentrated below an altitude of 3000 meters, we use the average temperature and atmospheric pressure of Hong Kong to compute the methane absorption cross-section and calculate the optical depth using the mean atmospheric density from 0 to 3000 meters. Although this approach might sacrifice some spectral simulation precision, the trade-off is considered acceptable due to the significant computational burden involved in generating the inversion dataset. The entire code for creating the inversion dataset is written in PyTorch and runs on an NVIDIA GeForce RTX 3060 GPU.
@@
[…]

### v3

@@ II-E Multi-task Learning I (MTL-01)
Considering that the loss function of Mask R-CNN measures instance segmentation and does not consider error propagation, we design a serial network (MTL-01), which takes the segmentation results of Mask R-CNN as input to ResNet-50 and calculates additional loss terms for back-propagation to assist in optimizing Mask R-CNN (Fig. 1). The ResNet-50 in the serial network is a pre-trained model, whose parameters are not updated. During the training process, the erroneous (over-)segmentation masks from Mask R-CNN are mainly categorized into two types. The first type is small false positive patches escaping from NMS(Non-Maximum Suppression) filtering[he2017mask], which were only found to slightly contribute to errors in subsequent emission rate estimation rather than main error source of the serial network. Simulation experiments indicate that the IME value of plumes is almost never below 300. Hence, we disregard false positive patches with IME values below 300 pixels, addressing this type of error using an object detection loss term. The second type of error consists of large false positive patches, often wrongly separated from overlapping plumes, causing substantial inaccuracies in emission rate estimation. To mitigate this, we introduce an additional loss term, $loss_{ER}$, for MTL-01. Although $loss_{ER}$ is still based on smoothL1, segmentation results with IME values below 300 pixels are excluded from its calculation. Missed masks are treated as having an emission rate estimate of 0 kg/h, and the loss is computed accordingly:
$loss_{ER}=\begin{cases}Smooth_{L1}(y_{i},\hat{y_{i}})&\text{if }y_{i}\in TP\\ Smooth_{L1}(0,\hat{y_{i}})&\text{if }y_{i}\in FP\end{cases}$ (16)
@@
The first multi-task learning model (MTL-01) is a sequential architecture combining a Mask R-CNN for plume segmentation and a ResNet-50 for estimating plume emission rates. The segmented plumes from Mask R-CNN are fed into ResNet-50 in real-time for emission rate loss calculation. The total loss for MTL-01 is the weighted sum of the individual losses of both networks. ’Multiplume’ describes a methane concentration map containing multiple plumes and noise. ResNet-50 is a 50-layer deep residual network. A Fully Convolutional Network (FCN) integrates semantic information from both low-level and high-level images via skip connections. In Mask R-CNN, ’heads’ are responsible for detecting objects and extracting masks. MLP stands for multi-layer perceptron. The second multi-task learning model (MTL-02) detects methane concentration and plume boundaries simultaneously. This model consists of a Mask R-CNN for plume segmentation and a U-Net for methane concentration inversion, both sharing a ResNet-50 feature extractor. While the Mask R-CNN’s decoding structure is unchanged, the U-Net decoding structure includes four skip connections, an ASPP module, and two 1x1 convolutional layers. ’Simulated EnMAP’ refers to EnMAP data augmented with simulated methane plume signals. The ASPP module comprises a 1x1 convolutional layer for channel reduction and multiple parallel atrous convolutional layers with adjustable dilation rates to capture features at different scales. ’Conv 1x1’ indicates a 1x1 convolution layer, ’Dilated Conv’ denotes an atrous convolution layer, and ’Pool’ refers to a pooling layer.
@@ II-F Multi-task Learning II (MTL-02)
The process of methane concentration inversion entails estimating pixel-specific values using hyperspectral images, whereas methane plume segmentation is concerned with identifying plume masks within the same sets of images. Given their structural similarity and logical relationship, using multi-task learning models might improve the accuracy of both methane concentration inversion and plume segmentation tasks. Hence, a multi-task learning framework that merges U-net and Mask R-CNN architectures has been devised.
In MTL-02 (Fig.1), the mask R-CNN network architecture remains unchanged, and the U-net decoding structure is connected to the mask R-CNN backbone network. Considering the output sizes of the stages in the ResNet-50 backbone network, we added an additional output to the first convolution block of ResNet-50 to ensure that the size of the U-net decoding result matches the input size, similar to LinkNet[zhou2018d]. It should be noted that the deepest output of ResNet-50 is not involved in the construction of the U-net decoding structure but only participates in the FCN[long2015fully] structure of Mask R-CNN. To address the potential underfitting of the U-net decoding structure due to the max-pooling layer situated between two outputs of ResNet-50, we introduced an Atrous Spatial Pyramid Pooling (ASPP) layer directly after the U-net decoder. This ASPP layer consists of two dilated convolutions with dilation rates of 2 and 4. ASPP[chen2017deeplab], which builds on dilated convolutions and SPP (Spatial Pyramid Pooling)[he2015spatial], addresses the challenge of balancing the reduction in feature map resolution with the need to capture a large receptive field for image feature extraction. ASPP leverages dilated convolutions with varying dilation rates to encapsulate the image context at multiple scales. In this approach, the input feature map is subjected to several dilated convolutions with different rates, which are then combined. To further expand the model’s receptive field, the resulting feature maps from these dilated convolutions are subjected to average pooling.
In the loss function for $MTL-02$, a traditional approach is employed to use a weighted average of the individual loss functions for multiple tasks[bruggemann2021exploring]:
@@
where $loss_{U{-}net}$ represents the loss function of U-net, $loss_{Mask\,R{-}CNN}$ denotes the loss function of Mask R-CNN, and $w_{1}$ along with $w_{2}$ are the respective weights for these loss functions.
To determine the weights for the MTL-02 loss function, we utilized the Dynamic Weighted Average (DWA) algorithm[liu2019end]. The main idea is that tasks differ in their difficulty and speed of learning. Instead of giving each task a separate learning rate, it is better to adjust the weights dynamically so that all tasks advance at a similar pace. In simple terms, tasks that show a faster decrease in loss receive smaller weights, while those with a slower decrease are allocated larger weights.
${w}_{k}(t-1)=\frac{L_{k}(t-1)}{L_{k}(t-2)}$ (19)
${\lambda}_{k}(t)=\frac{K\cdot e^{\frac{{w}_{k}(t-1)}{T}}}{\sum_{i=1}^{K}e^{\frac{{w}_{i}(t-1)}{T}}}$ (20)
where $K$ is the number of tasks, $\lambda_{k}(t)$ is the weight of the task k at the i-th iteration, and $L_{k}(t)$ is calculated as the average loss in each epoch over several iterations. When $T$ is a temperature, $T=1$ indicates that $\lambda$ is equal to the result of softmax; However, as $T$ increases significantly, $\lambda$ approaches 1, leading to equal loss weights for all tasks.
@@ III Experiments
III-A Research area and data
III-A1 Research area
Situated in the southern part of China, Hong Kong is predominantly marked by hilly terrain. The region, which is highly urbanized, experiences a humid and rainy climate, with an average annual precipitation exceeding 2000 millimeters. Within Hong Kong, there are two key landfills: West New Territories (WENT) landfill and North East New Territories (NENT) landfill. Since most of the PRISMA images were obtained from the WENT region, our generalization verification is mainly concentrated in this area. The Tuen Mun landfill (Fig. 2), which spans approximately 27 hectares, is located in the southwest of the Tuen Mun district near Shenzhen Bay. It is one of the limited waste disposal sites sanctioned by the government in the Hong Kong Special Administrative Region and primarily receives municipal solid waste from various areas, including Hong Kong Island, Kowloon and New Territories.
@@ III-A2 Hyperspectral Data
The PRISMA satellite was launched on 22 March 2019, and offers hyperspectral images of global coverage with a spatial resolution of 30 m. The spectral smile is less than 5 $nm$, the spectral resolution is better than 12 $nm$ in a spectral range of 400-2500 nm (VNIR and SWIR regions) [cogliati2021prisma], and datasets are open-access. We have collected several PRISMA images of WENT, which are: $20220102$, $20220119$, $20220914$, $20230217$, $20230301$, $20230307$, $20230416$. These raw images will be used to verify the generalizability of deep learning models[thorpe2013high][foote2020fast].
@@
The research site is the WENT landfill situated in the Tuen Mun District of Hong Kong SAR. The landfill is located on a hillside close to Shenzhen Bay, as indicated by the blue rectangle on the map.
@@
III-B Generation of diversified plumes in terms of emission rates and wind speeds
The simulated plumes were generated using Palm software[2001PALM][maronga2015parallelized],using the average atmospheric pressure at sea level in Hong Kong (0.96 hPa) and a temperature of 296K. A specialized smog mode was implemented, including a methane dispersion feature, to simulate the diffusion of methane point sources within the troposphere from 0 to 10,000 meters. We defined four gradients for the methane point source emission rate: 500kg/h, 1000kg/h, 1500kg/h, and 2000kg/h. The wind direction was set from west to east, with wind speed gradients of 1m/s, 2m/s, 3m/s, 4m/s, 5m/s, 6m/s, 7m/s, 8m/s, 9m/s, and 10m/s (Fig. 3). For each wind speed gradient, the methane point source emission rate gradients were simulated in sequence. The simulation period covered 0 to 2.5 hours, with snapshots taken at 30-second intervals from 1 to 2.5 hours, resulting in a total of 7200 plume snapshots. Among them, the simulated plume emission data with wind speeds of $2m/s$, and $10m/s$ were used to create the validation dataset, the simulated plume emission data with wind speeds of $1m/s$, and $9m/s$ were used to create the test dataset, while the remaining plume emission data were used to create the training dataset.
@@ III-B Generation of diversified plumes in terms of emission rates and wind speeds
This figure illustrates the methane plume at various emission rates and wind speeds, modeled with the Palm software. Each image was taken 180 minutes after the initiation of the emission, with simulation parameters specified in Section II.B.
@@ III-C Generation of multi-task learning datasets
By capturing an image of a plume and averaging the methane concentrations at various heights, a 2D representation of a single plume with an emission rate of $m$ kg/h is produced. Scaling all pixel values in this 2D plume by a factor of $a$ results in a new emission rate of $m\times a$ kg/h, which is used as the sample label in the emission rate dataset. To ensure plume diversity, we randomly generate a threshold between 0.05 and 0.10, and set pixels in the 2D plume below this threshold to zero. A rotation enhancement ranging from -170° to 170° is then applied to the 2D plume, creating the image counterpart for the emission rate dataset sample. Plume segmentation is treated as an instance segmentation task. Initially, we create a blank $256\times 256$ image filled with zeros. We then randomly select $N$ images from the emission rate dataset (where $N\leq 3$ and $N$ can be 0) and superimpose them onto the all-zero image to form a multi-plume image. Concurrently, we determine the boundaries of each plume and assign corresponding labels to the segmentation dataset. If $N=0$, the dataset will lack labels. Lastly, we randomly generate $256\times 256$ Gaussian white noise and add it to the multi-plume image[jongaramrungruang2022methanet][gil2002efficient] to create the image for the segmentation dataset. Additionally, the overlap ratio between any two plumes must not exceed 15
@@
The inversion data originate from multi-plume images created during the process of generating the segmentation dataset. These multi-plume images serve as labels for the inversion dataset samples. First, a multi-plume image is obtained and inspected for any emptiness. If found to be empty, a random EnMAP image is chosen as the sample image. On the other hand, if the multi-plume image is not empty, methane transmission spectra for each pixel value are constructed on a column-by-column basis. For columns that are entirely zeros, the transmission spectrum is defined as a $256\times 1\times 41$ matrix where all elements are set to 1. For columns with non-zero values, an optical depth matrix with dimensions $256\times 1\times K$ (where $K$ is the number of elements in the absorption cross-section dataset) is created using the approach explained in Section II.A. The negative logarithm of the pixel values in the optical depth matrix is then used to compute the original transmittance matrix.
First, the central wavelengths and half-bandwidths of the 41 EnMAP bands are determined to construct the spectral response function (SRF)[green1998spectral][2008Toward]. Then, the SRF is utilized for normalized convolution along the $K$ dimension of the initial transmittance matrix, resulting in a $256\times 1\times 41$ ’column’ transmission spectrum matrix. Following the processing of each column of the non-empty multi-plume image, all the transmission spectrum matrices are combined sequentially in their original column order to form a $256\times 256\times 41$ transmission spectrum matrix.
In the final step, the transmission spectrum matrix is merged with random EnMAP data through the Hadamard product to produce the image for the non-empty multi-plume. It’s important to note that, since the simulated methane plumes are mainly below 3000 meters in altitude, we use the average temperature and atmospheric pressure of Hong Kong to determine the methane absorption cross-section and calculate the optical depth based on the mean atmospheric density from 0 to 3000 meters. While this method may reduce some accuracy in spectral simulation, the compromise is worthwhile given the substantial computational load of creating the inversion dataset. The entire code for generating the inversion dataset is implemented in PyTorch and executed on an NVIDIA GeForce RTX 3060 GPU.
@@
The training dataset is composed of three types of samples: Emission rate sample, Segmentation sample, and Inversion sample. The training dataset consists of numerous examples of these types of samples. (a) A snapshot of a plume, simulated by Palm, is scaled (including the emission rate), and concentration values below a certain threshold are set to zero. A random rotation from -170° to 170° is applied to produce a single enhanced plume, which acts as the emission rate sample image. The scaled emission rate serves as its label. (b) Up to N ($N\leq 3$) plumes generated in step (a) are randomly merged (ensuring that overlap does not exceed 70%) to create a multi-plume image. Noise from mag1c inversion outcomes is added to this multi-plume image to generate the segmentation sample image, with the boundary of each plume noted as the label. (c) Following the method in (a), a transmission spectrum matrix that matches the EnMAP base map dimensions is generated based on the multi-plume image from (b). The simulated EnMAP image is then created by applying the Hadamard product of the transmission spectrum matrix and the EnMAP base map, which serves as the inversion sample. The multi-plume image serves as the label for the inversion sample.